# Lusztig sheaves, characteristic cycles and the Borel-Moore homology of Nakajima’s quiver varieties

### v1

@@ 6.4. Characteristic cycles for two-framed quivers
By Proposition 6.7, $\mathcal{Q}_{\nu,\omega^{1},\omega^{2}}$ is a full subcategory of $\mathcal{D}^{b}_{\mathbf{G}_{\mathbf{V}}\times\mathbf{G}_{\Omega}}(\mathbf{E}_{\mathbf{V}\oplus\mathbf{W}^{1}\oplus\mathbf{W}^{2},\Omega^{(2)}})$, hence $\mathcal{K}(\omega^{1},\omega^{2})_{-1}^{\psi}$ can be regarded as a $\mathbb{Z}$-submodule of $\mathcal{K}_{0}(\mathcal{D}^{b}_{\mathbf{G}_{\mathbf{V}}\times\mathbf{G}_{\Omega}}(\mathbf{E}_{\mathbf{V}\oplus\mathbf{W}^{1}\oplus\mathbf{W}^{2},\Omega^{(2)}})$. Compose the characteristic cycle map with $[\textrm{For}_{\mathbf{G}_{\mathbf{V}}}^{\mathbf{G}_{\mathbf{V}}\times\mathbf{G}_{\Omega}}]$, we obtain a morphism
@@
Given dimension vector $\nu=\nu^{\prime\prime}+ri$ and fix graded spaces $\mathbf{V}^{\prime\prime}\subseteq\mathbf{V}$ with dimension vectors $\nu^{\prime\prime}$ and $\nu$ respectively. Let $Y=\mathbf{E}_{\mathbf{V}^{\prime\prime},\mathbf{W}^{1}\oplus\mathbf{W}^{2},\Omega^{(2)}}$, $X^{\prime}=\mathbf{E}_{\mathbf{V}\oplus\mathbf{W}^{1}\oplus\mathbf{W}^{2},\Omega^{(2)}}$ and $V=F_{\mathbf{W}^{1}\oplus\mathbf{W}^{2}}$, $X=\mathbf{G}_{\mathbf{V}}\times^{P}Y$ and $W=\mathbf{G}_{\mathbf{V}}\times^{P}F_{\mathbf{W}^{1}\oplus\mathbf{W}^{2}}$. Then the induction correspondence of $\mathbf{Ind}$ in section 2.1.1 coincides with the diagram of the functor $\mathcal{F}^{(r)}_{i}$. Let $\bar{F}_{\mathbf{W}^{1}\oplus\mathbf{W}^{2}}$ be the subset of $T^{\ast}\mathbf{E}_{\mathbf{V}\oplus\mathbf{W}^{1}\oplus\mathbf{W}^{2},\Omega^{(2)}}\cong\mathbf{E}_{\mathbf{V},\mathbf{W},\Omega^{(1)}}$ consisting of $(x,y,z)$ such that $(x,y,z)(\mathbf{V}^{\prime\prime})\subseteq\mathbf{V}^{\prime\prime}$ and denote $\bar{F}_{\mathbf{W}^{1}\oplus\mathbf{W}^{2}}\cap\Pi_{\mathbf{V},\mathbf{W}^{1}\oplus\mathbf{W}^{2}}$ by $\bar{F}^{nil}_{\mathbf{W}^{1}\oplus\mathbf{W}^{2}}$. By [9, Lemma 9.4], we have $\mathbf{G}_{\mathbf{V}}\times^{P}\bar{F}_{\mathbf{W}^{1}\oplus\mathbf{W}^{2}}\cong T^{\ast}_{W}(X\times X^{\prime})$, and there is a commutative diagram,
@@
where $\Delta$ is the morphism in Theorem 4.6 and those ${\rm{CC}}^{s}$ are the morphisms in Theorem 6.16 and 6.23 respectively. Then after base change to $\mathbb{Q}$, $\delta$ is the unique isomorphism which satisfies the conditions in Conjecture 5.3.

### v3

@@ 6.4. Characteristic cycles for two-framed quivers
By Proposition 6.7, $\mathcal{Q}_{\nu,\omega^{1},\omega^{2}}$ is a full subcategory of $\mathcal{D}^{b}_{\mathbf{G}_{\mathbf{V}}\times\mathbb{T}}(\mathbf{E}_{\mathbf{V}\oplus\mathbf{W}^{1}\oplus\mathbf{W}^{2},\Omega^{(2)}})$, hence $\mathcal{K}(\omega^{1},\omega^{2})_{-1}^{\psi}$ can be regarded as a $\mathbb{Z}$-submodule of $\mathcal{K}_{0}(\mathcal{D}^{b}_{\mathbf{G}_{\mathbf{V}}\times\mathbb{T}}(\mathbf{E}_{\mathbf{V}\oplus\mathbf{W}^{1}\oplus\mathbf{W}^{2},\Omega^{(2)}})$. Compose the characteristic cycle map with $[\textrm{For}_{\mathbf{G}_{\mathbf{V}}}^{\mathbf{G}_{\mathbf{V}}\times\mathbb{T}}]$, we obtain a morphism
@@
Given dimension vector $\nu=\nu^{\prime\prime}+ri$ and fix graded spaces $\mathbf{V}^{\prime\prime}\subseteq\mathbf{V}$ with dimension vectors $\nu^{\prime\prime}$ and $\nu$ respectively. Let $Y=\mathbf{E}_{\mathbf{V}^{\prime\prime},\mathbf{W}^{1}\oplus\mathbf{W}^{2},\Omega^{(2)}}$, $X^{\prime}=\mathbf{E}_{\mathbf{V}\oplus\mathbf{W}^{1}\oplus\mathbf{W}^{2},\Omega^{(2)}}$ and $V=F_{\mathbf{W}^{1}\oplus\mathbf{W}^{2}}$, $X=\mathbf{G}_{\mathbf{V}}\times^{P}Y$ and $W=\mathbf{G}_{\mathbf{V}}\times^{P}F_{\mathbf{W}^{1}\oplus\mathbf{W}^{2}}$. Then the induction correspondence of $\mathbf{Ind}$ in section 2.1.1 coincides with the diagram of the functor $\mathcal{F}^{(r)}_{i}$. Let $\bar{F}_{\mathbf{W}^{1}\oplus\mathbf{W}^{2}}$ be the subset of $T^{\ast}\mathbf{E}_{\mathbf{V}\oplus\mathbf{W}^{1}\oplus\mathbf{W}^{2},\Omega^{(2)}}\cong\mathbf{E}_{\mathbf{V},\mathbf{W},\Omega^{(1)}}$ consisting of $(x,y,z)$ such that $(x,y,z)(\mathbf{V}^{\prime\prime})\subseteq\mathbf{V}^{\prime\prime}$ and denote $\bar{F}_{\mathbf{W}^{1}\oplus\mathbf{W}^{2}}\cap\Pi_{\mathbf{V},\mathbf{W}^{1}\oplus\mathbf{W}^{2}}$ by $\bar{F}^{nil}_{\mathbf{W}^{1}\oplus\mathbf{W}^{2}}$. By [14, Lemma 9.4], we have $\mathbf{G}_{\mathbf{V}}\times^{P}\bar{F}_{\mathbf{W}^{1}\oplus\mathbf{W}^{2}}\cong T^{\ast}_{W}(X\times X^{\prime})$, and there is a commutative diagram,
@@
where $\Delta$ is the morphism in Theorem 4.6 and those ${\rm{CC}}^{s}$ are the morphisms in Theorem 6.16 and 6.24 respectively. Then after base change to $\mathbb{Q}$, $\delta$ is the unique isomorphism which satisfies the conditions in Conjecture 5.3.
@@
Remark 6.27.
Indeed, after using the same analysis of Section 12 in [27], one can deduce that the homology groups $\mathbf{H}(\mathfrak{L}(\omega),\mathbb{Z})$ and $\mathbf{H}(\tilde{\mathfrak{Z}},\mathbb{Z})$ have ${}_{\mathbb{Z}}\mathbf{U}(\mathfrak{g})$-module structures.